OpenStudy (anonymous):

Write as a power of 7 7^5 x 7^3

Nnesha (nnesha):

when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

OpenStudy (anonymous):

Thank you Nnesha

Nnesha (nnesha):

np :=) so what did you get ?

OpenStudy (ahvei):

would you like a direct answer???

OpenStudy (ahvei):

or i can work it out a different way for you

Nnesha (nnesha):

doesn't matter if they like it or not its just against the rules.

OpenStudy (anonymous):

49^15? Its probably wrong xD

Nnesha (nnesha):

base would stay the same just add their exponents!

OpenStudy (ahvei):

its suppose to be multiplied and thats still wrong

OpenStudy (anonymous):

7x7^15?

OpenStudy (anonymous):

Realise that now, got confused.

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\] \(\color{blue}{\text{End of Quote}}\) ADD not multiply

OpenStudy (anonymous):

7x7^8

Nnesha (nnesha):